MICROWAVE SPECTROSCOPY AND POTENTIAL ENERGY SURFACES OF Ne-OH and Kr-OH

\underline{YOSHIHIRO SUMIYOSHI}, IPPEI FUNAHARA, KAZUYA SATO, YASUHIRO OHSHIMA,^{a} and YASUKI ENDO, Department of Basic Science, Graduate School of Arts and Sciences, The University of Tokyo, Komaba, Meguro-ku, Tokyo 153-8902, Japan.

Three-dimensional potential energy surfaces for Ne-OH/OD and Kr-OH/OD have been determined from the data by Fourier-transform microwave spectroscopy with fine and hyperfine structures. For Ne-OH, data of rovibrational transitions between the levels with P = 0.5 and 1.5 have been utilized as well as those of the pure rotational transitions. Simultaneous least-squares analyses utilizing the data of the two isotopollogues were performed, where all the freedom of motions for an atom-diatom system in the Jacobi coordinate, R, θ, and r, were considered. The initial values for the least-squares analyses were obtained by high-level ab initio calculations at the UCCSD(T)-F12 theory with the basis set of aug-cc-pV5Z for Ne-OH and aug-cc-pVQZ for Kr-OH. Dependences of the intermolecular interaction energy on the bond length of the OH monomer of the two complexes are compared with that of Ar-OH ^{b}, where the interaction potential of Ar-OH has been redetermined usging the new initial potential parameters by UCCSD(T)-F12/aug-cc-pV5Z.

aPresent address: Institute for Molecular Science, Myodaiji, Okazaki 444-8585, Japan

bY. Sumiyoshi, I. Funahara, K. Sato, Y. Ohshima, and Y. Endo, J. Chem. Phys. \underline{125}, 124307 2006.